Mathematics 134 – Intensive Calculus for Science 2 Discussion 2 –“Integral Derby” February 6, 2002

Directions

• First 20 minutes: Each group makes up three integrals. Any formula that can be integrated by the rules we have discussed in class so far – that is: the power, exponential, sin, cos, sum and constant multiple rules, and substitution – possibly after some preliminary algebraic rearrangement is OK.

• Prepare an answer key, giving the answers for your integrals.

• Make an identical copy of your list of functions for each member of your partner group, and one answer key.

• Second 15 minutes: Each pair of partner groups exchanges functions. Find the general antiderivatives of the functions you are given, working as a group. Prepare one set of solutions.

• Final 15 minutes: (The moment of truth!) Partner groups combine and compare both sets of solutions with the answer keys to compute scores according to the rules below.

Scoring Rules

1) Each correct integral from your partner group you compute is worth 5 points.

2) If your group “stumped” your partner group, and you computed your integral correctly in the answer key, your group gets 3 bonus points.

3) For each incorrect solution in your group’s answer key, though, you lose 5 points.

4) Answers should still be counted as correct if they are equivalent to the answer key but in a different form (for example: written with radicals rather than fractional powers).

5) Best total group score “wins” the Derby. (But of course, everyone wins in the end!)

6) The referee for any disputed call about whether a function is legal or whether an answer is correct is Pr. Little. It’s OK to ask the referee to make a ruling before you give your problems to your partner group, if you are in doubt.